OpenStudy (butterflyhope):

Check *will medal*: 1 5/6 - 1 1/8 The LCM of 6 and 8 is 24. So 6 x 4 = 24 and 5 x 4 = 20 = 20/24. And then 8 x 3 = 24 and 1 x 3 = 3 = 3/24. 1 20/24 - 1 3/24 = 17/24. Thanks c:

OpenStudy (animal_lover36):

Thats correct.

OpenStudy (anonymous):

1 5/6 1 1/8 ----- temporarily take out the whole numbers...just deal with fractions 5/6 = 20/24 1/8 = 3/24 now put the whole numbers back in and subtract 1 5/6 - 1 1/8 = 1 20/24 - 1 3/24 = 17/24 yep...you got it

OpenStudy (butterflyhope):

Thanks c: